OpenStudy (anonymous):

Please Check My Work! Medal/Fan The grades on the last history exam had a mean of 75%. Assume the population of grades on history exams is known to be normally distributed with a standard deviation of 7. What percent of students earn a score between 70% and 82%? A. 0.6285 B. 0.6101 C. 0.6024 D. 0.5892

OpenStudy (anonymous):

I said C for this one

OpenStudy (anonymous):

@ParthKohli can u help with stats?

OpenStudy (tkhunny):

Have you considered calculating the Z-scores?

OpenStudy (tkhunny):

\(\dfrac{70-75}{7} = -5/7 = -0.714285\) You do the other one.

OpenStudy (anonymous):

okay, so is it going to be 70-82 ------- 7 ?

OpenStudy (anonymous):

so how do you know which numbers to use?

OpenStudy (tkhunny):

No good. Use the mean. We need to know how many standard deviations above the mean (75). (82-75)/7 = ??

OpenStudy (anonymous):

oh, the mean is 75 .... OH!!!! okay, so its 7/7 which is 1

OpenStudy (tkhunny):

No, we must determine how much Standard Normal probability lies between Z = -0.714285 and Z = +1. This can be done many ways. You probably have a calculator that will do it.

OpenStudy (anonymous):

ok, so is it averaging? I'm really sorry, i'm kinda new to this thing

OpenStudy (tkhunny):

You should have some sort of cumulative distribution table or calculator function. In MS Excel 2013 it is this =NORM.DIST(-0.714285,0,1,1) = 0.237525. This suggests that 23.7525% of the probability of the Standard Normal Distribution lies to the left of -0.714285 standard deviations from the mean. In a table, and there are many... https://www.stat.tamu.edu/~lzhou/stat302/standardnormaltable.pdf We read the values: Z = -0.71 ==> 0.23885 and Z = -0.72 ==> 0.23576. Interpolating might give Z = -0.714285 ==> 0.23753. This is very close to the value MS Excel gave us. You have to find a way that you understand and that is available to you.

OpenStudy (anonymous):

ok, thank you for your help :)

OpenStudy (tkhunny):

After that, you must do the same thing for Z = +1. Tables and calculators are different. You have to understand your output. In my examples, I calculated probability to the left of the Z-Score. Another table may be to the right. In this case, the answer is: Using MS Excel P(Z < 1) - P(Z < -0.714285) = 0.841345 - 0.237525 = 0.603819 Using the table and interpolation 0.603813 Make sure you understand both the problem statement and the tools. Either one lacking can make it all go horribly wrong. :-)

OpenStudy (anonymous):

ok :) I'll look into that :D